A firm in monopolistic competition has the firm demand curve: P = 60 - 2Q. The Total Cost equation is TC = 40 + Q2

How much deadweight loss is created by the firm?
Answer #1

Note that MR = 60 - 4Q and MC = 2Q. This gives profit maximizing quantity at MR = MC or 60 - 4Q = 2Q. This gives Q = 10 units and P = $40. Now P = MC gives 60 - 2Q = 2Q or Q = 60/4 = 15 units and price P = 0.5

DWL = 0.5*(Monopolistically competitive price - MR = MC)*(Competitive quantity - Monopolistically competitive quantity)

= 0.5*(40 - 20)*(15 - 10)

= $50

Hence the deadweight loss is $50.